# VULCA-Bench: A Multicultural Vision-Language Benchmark for Evaluating Cultural Understanding

### v1

@@ 4 Dataset Statistics
Vulca-Bench comprises 7,410 matched image-critique pairs drawn from 11,964 total images across 8 cultural traditions, annotated with 225 culture-specific dimensions. Table 5 summarises key metrics: average Chinese critique length is 450 characters (English: 280 words), with 100% bilingual completeness, zero duplicates, >99% metadata verification, and 98% cultural-fact accuracy. All pairs meet the $\geq$70% dimension coverage threshold. Western (54.5%) and Chinese (27.6%) dominate, while minority cultures maintain rigorous verification. Audit levels range from full manual review (Korean) to stratified sampling (Indian).
@@
Figure 2 reveals layer-wise coverage patterns: L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), while L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding. Chinese art shows highest L5 coverage (78%), consistent with its rich philosophical aesthetics tradition (qiyun, yijing). Importantly, each culture’s critiques are authored by native expert annotators from that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapanese
Figure 2: Layer-wise dimension coverage. L1–L2 achieve $\geq$94%; L3–L5 decline to 72–89%, reflecting interpretive complexity.
5 Data Quality Validation
We validate Vulca-Bench through a three-phase audit protocol. Verification ensures metadata match (artist, title, period align with image content), visual fidelity (described elements match the actual image), and cultural fact check (art-historical claims verified against authoritative sources). Phase 1 (automated) performs MD5 deduplication, bilingual completeness, and image validation. Phase 2 (stratified) constitutes the 450 human-scored subset. KR: 107 (full set), JP/IS/IN: 50 each, CN/WE: 100 each. Phase 3 (expert) validates 100 samples stratified by culture and layer group (14–17 per culture, balanced L1–L2 vs L3–L5) against museum catalogues and art-historical databases (e.g., Grove Art Online, Oxford Art Online). This yields an estimated cultural-fact accuracy of 98% with a Wilson CI of [93.0%, 99.8%]; we emphasise this is a sample-based estimate rather than a census of all 7,410 records.
Reproducibility: Stratification uses a fixed random seed (42) with proportional allocation by culture size. Per-sample audit outcomes (pass/fail/corrected) are logged in audit_log_p3.csv (supplementary), and full reviewer assignments are documented in Appendix E.
Table 6 reports quality metrics with Wilson confidence intervals: 100% bilingual completeness, 0% duplicates, 100% image availability (100-sample audit), and 100% dimension coverage ($\geq$70% threshold). To address sample-size imbalance concerns, we provide a balanced subset (N=48 per culture $\times$ 7, total 336 pairs). Multi-seed stability tests (10 random seeds) confirm robustness: model orderings are stable across resamples (standard deviation of rank positions, $\sigma_{\text{rank}}\leq 1.5$). Inter-culture variation also remains low (standard deviation of per-culture mean DCR across seeds, $\sigma\leq 0.05$). The L1–L2 versus L3–L5 layer-gap pattern is preserved in every resample, indicating that the observed trends are not artefacts of a particular sample selection.
Metric Val 95% CI
Bilingual complete 100% [99.9%, 100%]
Duplicate rate 0% [0%, 0.1%]
Image available 100% [96.3%, 100%]
Dim $\geq$70% 100% [99.9%, 100%]
Table 6: Quality metrics with Wilson CI. 100-sample audit.
Vulca-Bench supports multiple research applications. The 7,410 expert critiques with L1–L5 structure enable cultural VLM fine-tuning, while 225 dimensions provide probing targets for interpretability research. The balanced subset (N=336 for pilot, 48/culture $\times$ 7; N=384 available for full evaluation with 8 cultures) enables equal-weighted cultural fairness probing, and bilingual critiques support cultural knowledge retrieval and RAG augmentation. [^1] Hermitage excluded from pilot due to late inclusion; full 8-culture evaluation (N=384) available in released dataset.
@@ 6 Pilot Evaluation
To demonstrate Vulca-Bench’s utility as a cultural VLM benchmark, we evaluate 5 representative VLMs on a stratified sample. This pilot evaluation establishes dataset utility and diagnostic capability only; it does not constitute a definitive evaluation framework. We use Dimension Coverage Rate (DCR) as a dataset-level diagnostic check, verifying that VLM outputs engage with culture-specific concepts, rather than as a leaderboard metric for model ranking.
@@ 6.1 Setup
We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash (temperature=0.7, max_tokens=2048; full configuration in supplementary). We use Dimension Coverage Rate (DCR) as a diagnostic to test whether model-generated critiques engage with culture-specific concepts. For a critique $c$ from culture $k$:
@@
where $\mathcal{D}_{k}$ is the culture-specific dimension set (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$) and $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ denotes the subset of dimensions detected via keyword matching. Each record’s covered_dimensions field (Section 3) provides the gold set, namely dimension IDs explicitly labeled by annotators during critique creation. All models generate English critiques; DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”).
DCR is intended as a coarse diagnostic rather than a leaderboard metric. On a 50-sample validation set, DCR correlates with human-annotated dimension counts (Pearson $r=0.82$). Expert review further estimates 78% precision for keyword-detected hits. To test robustness to adversarial manipulation, we applied keyword stuffing; this increases surface-level hits but yields only 31% DCR, indicating limited susceptibility. For semantic validation, we report agreement rates between keyword detections and two independent checks at the sample level: (i) embedding-based similarity (cosine $\geq 0.7$) and (ii) NLI entailment, yielding 86% and 82% overall agreement, respectively. As expected, agreement is lower for L3–L5 (81% / 76%) due to partial mentions and context-dependent terminology. Full protocol details (thresholds, dictionaries, and length-controlled analyses) appear in Appendix D.
6.2 Dimension Coverage Results
@@
Table 7: Layer-Gap Diagnostic results. $\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit. All models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5). These values should be interpreted as a reference rather than state-of-the-art results because DCR is a coarse proxy and the pilot uses a limited prompt configuration. N=336 balanced-pilot subset (7 cultures); bootstrap 95% CI half-width $\approx$ 4.8 pp.
All models exhibit a 25–40 percentage-point layer gap ($\Delta_{\text{L}}=\text{DCR}_{\text{L1--L2}}-\text{DCR}_{\text{L3--L5}}$), confirming Vulca-Bench captures cultural understanding beyond visual perception. The 14-point DCR spread (72.4%–58.2%) demonstrates discriminative power across model capabilities. Critically, model orderings on the balanced-pilot subset (N=336, 7 cultures) correlate strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C). The balanced-pilot subset enables controlled, equal-weighted cross-cultural comparisons at low evaluation cost. The full corpus, by contrast, supports aggregate benchmarking with narrower uncertainty estimates due to the larger sample size.
@@ 6.3 Implications
These results confirm our hypothesis that cultural understanding is fundamentally hierarchical. The consistent 25–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability tier from visual perception. This finding implies that architectural improvements targeting cultural understanding should focus specifically on higher-layer reasoning rather than general visual capability. We release and validate DCR (78% precision; resistant to adversarial manipulation) as a reproducible, dataset-level diagnostic check signal for dimension coverage.
@@ 6.4 Error Analysis: VLM Failures on L3–L5
We identify three recurring error patterns (detailed examples in Appendix H): (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c. vanitas to 16th-c. pronkstilleven); (3) Cultural conflation—confusing distinct traditions despite clear stylistic markers (e.g., Persian vs. Mughal miniatures). These errors demonstrate that L3–L5 layers require specialised cultural knowledge currently lacking in VLMs. Vulca-Bench’s 195 dimensions provide diagnostic precision to categorise such failures systematically.
@@ 6.5 Few-Shot Learning with Expert Critiques
We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs). We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars (Table 8).
Contrary to expectations, few-shot prompting yields decreased performance across all tested VLMs. DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$). GPT-4o follows with $15.5\%$ decline ($p<0.001$). Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
We hypothesise three contributing factors: (1) attention dilution—longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting—exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity—exposure to expert critique structures may constrain flexibility. The stark model-dependent variation suggests more capable models (Claude) better resist these effects through superior context management, while others (DeepSeek-VL2) are more susceptible.
@@
7 Release and Licensing
The complete dataset is available under CC BY 4.0 in the supplementary materials. The release package includes: metadata JSON (7,410 pairs with bilingual critiques and dimension annotations), all images (optimised to $\leq$3.75MB), 225 dimension definitions, synonym dictionaries, evaluation scripts, and annotation guidelines. We release the full evaluation package and provide cached model outputs to support reproducibility. Using cached responses, the pilot experiments can be replicated at under USD 20 in API costs at the time of writing (see reproduce_pilot.py for the exact configuration). All artworks are from public museum collections (public domain or CC0); each record includes source_institution and accession_number for provenance (see Appendix I for culture distribution visualization).
@@ 8 Conclusion
We present Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, operationalised through a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions. The dataset addresses critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
Vulca-Bench enables systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception. The complete dataset (images, metadata, critiques, dimension definitions, and evaluation/annotation tools) is available under CC BY 4.0 in the supplementary materials. We emphasise that DCR serves as a coarse-grained diagnostic rather than a leaderboard metric.
Vulca-Bench enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset. The dataset complements, not replaces, human cultural expertise.
@@
Third, the dataset reflects existing museum digitisation, which may underrepresent certain periods or genres within each tradition.
@@
Finally, DCR serves as a coarse diagnostic rather than a precision instrument, designed for dataset utility validation rather than definitive model ranking.
@@ Appendix A Dimension Definitions
Full dimension definitions for each culture are provided in the supplementary materials. Example dimensions:
@@ Appendix B Dataset Variants
Eval 7,214 7 195 VLM evaluation (excl. Hermitage)
@@
Balanced 384 8 225 Fairness analysis (48/culture)
Balanced-Pilot 336 7 195 Pilot evaluation (excl. Hermitage)
Table 9: Dataset variants. C=cultures, D=dimensions. 8-culture: Full, Balanced (include Hermitage). 7-culture: Eval, Balanced-Pilot (exclude Hermitage). 6-culture: Gold, Human (exclude Mural, Hermitage due to insufficient gold annotations). Hash b8a34e5f* verifies single source.
@@ Appendix F Sample Critiques
Appendix G Open-Source Embedding Replication
To ensure reproducibility without proprietary API dependencies, we replicated the semantic alignment analysis (Section 6) using the open-source BGE-large-en-v1.5 model. [^2] Xiao et al., 2023. C-Pack: Packaged Resources To Advance General Chinese Embedding. https://arxiv.org/abs/2309.07597
Layer OpenAI BGE $\Delta$
L1–L2 94% 92% –2%
L3–L5 81% 79% –2%
Overall 86% 84% –2%
Table 11: Semantic alignment comparison: OpenAI text-embedding-3-small vs BGE-large-en-v1.5 (open-source). Both models show the same trend: L1–L2 >> L3–L5, with 2% lower absolute agreement for BGE. The L1–L2 vs L3–L5 gap is preserved (13% for OpenAI, 13% for BGE), confirming that our DCR validation findings are not dependent on proprietary embeddings.
Reproducibility: BGE-large-en-v1.5 is available on Hugging Face (BAAI/bge-large-en-v1.5) under MIT license. We release the replication script (eval/semantic_align_bge.py) with the dataset.
Appendix H Error Pattern Examples
Surface-level terminology.
GPT-4o analysing a Song Dynasty landscape scroll produces: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations. Expert correction identifies three visual strategies: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs. dian), and strategic negative space (liubai).
Historical anachronism.
Cultural conflation.
Gemini-2.5-Pro identifies a Safavid Persian miniature as “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
Appendix I Culture Distribution Visualization

### v2

@@ 4 Dataset Statistics
Vulca-Bench comprises 7,410 matched image-critique pairs drawn from 11,964 total images across 8 cultural traditions, annotated with 225 culture-specific dimensions. Table 5 summarises key metrics. The average Chinese critique length is 450 characters, while English critiques have an average of 280 words. Bilingual coverage was complete with all samples having EN-CN critiques, and all duplicates were excluded. All pairs meet the $\geq$70% dimension coverage threshold. Western (54.5%) and Chinese (27.6%) dominate, while minority cultures maintain rigorous verification. Audit levels range from full manual review (Korean) to stratified sampling (Indian).
@@
Figure 2 reveals layer-wise coverage patterns. While L1–L2 (visual/technical) achieve near-complete coverage ($\geq$94%), L3–L5 (cultural/philosophical) show progressive decline to 72–89%, reflecting the interpretive complexity of deeper cultural understanding. Chinese art shows highest L5 coverage (78%), consistent with its rich philosophical aesthetics tradition (qiyun, yijing). Importantly, each culture’s critiques are authored by annotators with specialized expertise in that tradition (Table 3), ensuring coverage patterns reflect indigenous aesthetic vocabulary depth rather than external cultural bias.
L1L2L3L4L5707080809090100100LayerCoverage (%)ChineseWesternJapaneseKoreanIslamicIndianMuralHermitage
Figure 2: Layer-wise dimension coverage across all right cultures. L1–L2 achieve $\geq$93%; L3–L5 decline to 70–89%, reflecting interpretive complexity.
@@ 5 Pilot Evaluation
Gemini-2.5-ProQwen3-VL-235BGPT-4oClaude-Sonnet-4.5GLM-4V-Flash02020404060608080100100Coverage (%)(a) Layer-wise CoverageL1–L2 (Visual/Technical)L3–L5 (Cultural/Philosophical)DeepSeek-VL2GPT-4oClaude-OpusClaude-Sonnet05050100100Coverage (%)(b) Few-shot Prompting Results0-shot1-shot3-shot
Figure 3: VLM diagnosis reveals systematic cultural understanding gaps. (a) Layer-wise coverage: all models exhibit 31–40 percentage-point gaps between L1–L2 and L3–L5, confirming cultural/philosophical layers are harder than visual/technical analysis. (b) Few-shot prompting results: DeepSeek-VL2 and GPT-4o show significant degradation from 0-shot to 3-shot; Claude models remain robust.
To demonstrate Vulca-Bench’s utility as a cultural VLM benchmark, we evaluate 5 representative VLMs on a stratified sample. This pilot evaluation is aimed at showcasing dataset utility and does not represent a definitive benchmark to compare against. For that purpose, we use Dimension Coverage Rate (DCR) to verify that VLM outputs engage with culture-specific concepts and to highlight the usefulness of Vulca-Bench, specifically focusing on cultural coverage rather than on a complete assessment of the VLM-generated critiques.
@@ 5.1 Setup
We sample 48 pairs per culture (N=336 total) and evaluate 5 VLMs: GPT-4o, Claude-Sonnet-4.5, Gemini-2.5-Pro, Qwen3-VL-235B, and GLM-4V-Flash. Temperature was set to 0.7 and max_tokens to 2048. We then use DCR as a measure to test whether model-generated critiques engage with culture-specific concepts. For a critique $c$ from culture $k$:
@@
where $\mathcal{D}_{k}$ is the culture-specific dimension set (e.g., $|\mathcal{D}_{\text{Chinese}}|=30$) and $\mathcal{D}_{k}^{c}\subseteq\mathcal{D}_{k}$ denotes the subset of dimensions detected via keyword matching. Each record’s covered_dimensions field provides the gold set, namely dimension IDs explicitly labelled by annotators during critique elaboration. All our models generate critiques in English language, then DCR is computed via keyword matching with culture-specific synonyms (e.g., “chiaroscuro” $\approx$ “light-dark contrast”). A complete description of the DCR validation protocol can be found in Appendix D.
5.2 DCR Results
@@
Table 6: DCR results. $\Delta_{L}$ = L1–L2 - L3–L5 measures cultural depth deficit. Results on N=336 balanced-pilot subset (7 cultures).
As observed in Table 6 (Figure 3, left), all models show a 31–40 percentage-point drop from visual perception (L1–L2) to cultural interpretation (L3–L5) i.e., $\Delta_{\text{L}}$, confirming that Vulca-Bench captures cultural understanding beyond just visual perception. The 14-point DCR spread (72.4%–58.2%) also highlights cultural-coverage differences across model capabilities. Interestingly, model performance on the balanced-pilot subset (N=336, 7 cultures) also correlates strongly with the full corpus (Spearman $\rho=0.94$ [0.87, 0.98]; Appendix C). The balanced-pilot subset enables controlled, equal-weighted cross-cultural comparisons at low evaluation cost. On the other hand, the full corpus supports aggregated benchmarking with narrower uncertainty estimates due to the larger sample size.
@@ 5.3 Implications
These results confirm our initial hypothesis that cultural understanding is fundamentally hierarchical. The consistent 31–40 percentage-point layer gap across all five models suggests that L3–L5 reasoning represents a qualitatively different capability from visual perception. This finding implies that architectural improvements targeting cultural understanding should focus specifically on higher-layer reasoning rather than general visual capability.
@@ 5.4 Error Analysis: VLM Failures on L3–L5
We identify three recurring error patterns: (1) Surface-level terminology—VLMs cite cultural terms (e.g., qiyun shengdong) without explaining concrete visual manifestations; (2) Historical anachronism—applying later artistic conventions to earlier works (e.g., 17th-c. vanitas to 16th-c. pronkstilleven); (3) Cultural conflation—confusing distinct traditions despite clear stylistic markers (e.g., Persian vs. Mughal miniatures).
Surface-level terminology.
GPT-4o analysing a Song Dynasty landscape scroll generates the following: “The painting embodies qiyun shengdong (spirit resonance)… The brushwork reflects the Six Principles, particularly qiyun shengdong.” The model correctly identifies the term but repeats it without explaining concrete visual manifestations.
Expert comment identifies three visual strategies instead: atmospheric perspective via layered mist, rhythmic brushwork variation (cun vs. dian), and strategic negative space (liubai).
Historical anachronism.
Cultural conflation.
Gemini-2.5-Pro identifies a Safavid Persian miniature as: “Mughal miniature with Rajput color palette and Hindu mythological elements.” Despite clear markers—Nasta’liq inscriptions (not Devanagari), Bihzad-influenced proportions, Persian shikar iconography without Hindu deities—the VLM conflates distinct Islamic and Hindu traditions.
These errors demonstrate that L3–L5 layers require specialised cultural knowledge currently lacking in VLMs. Vulca-Bench’s 225 dimensions provide diagnostic precision to categorise such failures systematically.
@@ 5.5 Few-Shot Learning with Expert Critiques
We evaluate whether Vulca-Bench’s expert critiques can serve as few-shot exemplars to improve L3–L5 coverage through cross-model validation (NN=1,028 across 4 VLMs). We compare zero-shot, 1-shot, and 3-shot prompting with DeepSeek-VL2, GPT-4o, Claude-Opus, and Claude-Sonnet, prepending $k$ culture-matched expert critiques as in-context exemplars.
Contrary to expectations, as observed in Table 7 and Figure 3 (right), few-shot prompting yields decreased performance across all tested VLMs. DeepSeek-VL2 shows the largest degradation: from 0.985 DCR at zero-shot to 0.578 at 3-shot ($41.3\%$, $p<0.001$). GPT-4o follows with $15.5\%$ decline ($p<0.001$). Interestingly, Claude models demonstrate remarkable robustness with $2\%$ degradation that is not statistically significant.
We hypothesise three contributing factors: (1) attention dilution, longer context from prepended exemplars may disperse focus from the target artwork; (2) style overfitting, exemplars may bias models toward mimicking surface-level formatting rather than genuine cultural reasoning; (3) template rigidity, exposure to expert critique structures may constrain flexibility. The noticeable model-dependent variation suggests more capable models (i.e., Claude) better resist these effects through superior context management, while others (e.g., DeepSeek-VL2) are more susceptible.
@@ 6 Conclusion
We presented Vulca-Bench, a multicultural art critique benchmark with 7,410 matched image-critique pairs across 8 cultural traditions, annotated considering a five-layer cultural understanding framework (L1–L5) with 225 culture-specific dimensions. The dataset addresses three critical gaps in existing VLM benchmarks: hierarchical cultural understanding, non-Western aesthetic frameworks, and expert-quality critique data.
Vulca-Bench, therefore enables a systematic evaluation of VLMs’ cultural capabilities beyond surface-level visual perception, focusing on critical interpretive aspects such as historical context or philosophical aesthetics. Vulca-Bench also enables research into culturally aware AI systems, including training data for culturally grounded VLMs, probing targets for interpretability research, and evaluation resources for cultural fairness probing via the balanced subset. We also report first-of-its-kind results on multi-layered cultural coverage, highlighting that while state-of-the-art VLMs tend to correctly assess visual and technical aspects, they struggle with deeper concepts such as cultural symbolism, historical context and philosophical aesthetics.
@@
Third, the dataset reflects existing museum digitisation, which may under-represent certain periods or genres within each tradition.
@@
Finally, DCR serves as a coarse diagnostic rather than a thorough evaluation, designed for dataset utility validation rather than definitive model ranking.
@@ Appendix A Dimension Definitions
Full dimension definitions for each culture are available at https://github.com/yha9806/VULCA-Bench. Example dimensions:
@@ Appendix B Dataset Variants
Eval 7,214 7 195 VLM eval (excl. Hermitage)
@@
Balanced 384 8 225 Fairness (48/culture)
Balanced-Pilot 336 7 195 Pilot (excl. Hermitage)
Table 8: Dataset variants. C=cultures, D=dimensions. 8-culture: Full, Balanced (incl. Hermitage). 7-culture: Eval, Balanced-Pilot (excl. Hermitage). 6-culture: Gold, Human (excl. Mural, Hermitage).
@@ Appendix F Sample Critiques
Appendix G Culture Distribution Visualization